# Fourier Analysis of Finite Difference Schemes for the Helmholtz Equation: Sharp Estimates and Relative Errors

### v1

@@ 3.3 Aliasing error of the classical 3-point centred scheme
$:=\sum_{(\xi/\pi)=1}^{N-1}\left(\frac{\xi}{\lambda}-\frac{\xi}{\lambda^{h}}\right)\hat{f}(\xi)\cos(\xi x)-\sum_{(\xi/\pi)=1}^{N-1}\frac{\xi}{\lambda^{h}}\sum_{s=\pm 1}s\sum_{m=1}^{\infty}\hat{f}(s\xi+2mN\pi)\cos(\xi x).$
@@
Suppose $k>\pi$, $10<\frac{kh}{2}\leq C_{\mu}<$ 1, and there is a sequence of $h=\frac{1}{N}$ going to zero such that
$\tilde{\sigma}_{k}:=\min_{\xi,h}\left|k-\frac{2}{h}\sin\frac{\xi h}{2}\right|>0\;\text{over $\xi\in\{1,\ldots,N-1\}\pi$ and the sequence of $h$ going to zero}.$
@@
because by the mean value theorem $\tilde{\theta}<\theta_{+}$. Note that the derivative of the numerator of $\phi^{\prime}(\theta)$ is $-2\theta\cos(2\theta)$. So if $\theta_{+}<\frac{\pi}{4}$, then $\phi^{\prime}<0$ on $[\theta_{+},\frac{\pi}{4}]$. Therefore, no matter $\theta_{+}<\frac{\pi}{4}$ or not, any stationary point $\theta_{c}$ of $\phi$ in $[\theta_{+},\theta_{\max}]$ must satisfy $\theta_{c}\geq\frac{\pi}{4}$. It follows that
@@
By Lemma 8 the second part $(E_{2}^{h})^{\prime}$ of $(e_{1}^{h})^{\prime}$ in (21) is controlled, similarly to $(e_{2}^{h})^{\prime}$ in (4), by the grid under-resolved source $f_{\mathrm{high}}$, and behaves as $O(h^{p})$ when $k>\pi$, $\frac{kh}{2}\leq C_{\mu}<1$ and $\tilde{\sigma}_{k}>0$. Moreover, the $O(h^{p})$ errors $|e_{2}^{h}|_{1}$ and $|E_{2}^{h}|_{1}$ do not depend directly on $k$ (but $\tilde{\sigma}_{k}$ for $|E_{2}^{h}|_{1}$).
@@
Suppose $k>\pi$, $10<\frac{kh}{2}\leq C_{\mu}<$ 1, and there is a sequence of $h=\frac{1}{N}$ going to zero such that
$\tilde{\sigma}_{k}:=\min_{\xi,h}\left|k-\frac{2}{h}\sin\frac{\xi h}{2}\right|>0\;\text{over $\xi\in\{1,\ldots,N-1\}\pi$ and the sequence of $h$ going to zero}.$
@@
$\leq$  $\;2\max_{\xi\in\{1,..N-1\}\pi}\frac{1}{|\lambda^{h}|^{2}}\sum_{(\xi/\pi)=1}^{N-1}\left[\left|\sum_{m=1}^{\infty}\hat{f}(\xi+2mN\pi)\right|^{2}+\left|\sum_{m=1}^{\infty}\hat{f}(\xi+2mN\pi)\right|^{2}\right].$
@@ 3.4 Absolute error estimates of the classical 3-point centred scheme
$\tilde{\sigma}_{k}:=\min_{\xi,h}\left|k-\frac{2}{h}\sin\frac{\xi h}{2}\right|>0\;\text{over $\xi\in\{1,\ldots,N-1\}\pi$ and the sequence of $h$ going to zero}.$
@@
where $\psi$ is given in (29), $\psi(k_{\pm}^{h})$ can be removed if $k_{\pm}^{h}\geq\xi_{\max}$, and $k_{+}$ (resp. $k_{-}^{h}$) coalesces with $k_{+}^{h}$ (resp. $k_{-}$) when $\pi\mathbb{N}\cap(k,k^{h})=\emptyset$. Moreover, we have
@@
[…]
@@
Note that $\pi\mathbb{N}\cap(k,k^{h})\neq\emptyset$ if and only if $k_{+}<k^{h}$. Using $\mu^{3}/6<\arcsin\mu-\mu<(\pi-2)\mu^{3}/2$ for $10<\mu<$ 1, we find a sufficient condition for $k_{+}\geq k^{h}$ is that $k^{3}h^{2}\leq 8\sigma_{+}/(\pi-2)$ and a sufficient condition for $k_{+}<k^{h}$ is that $k^{3}h^{2}>24\sigma_{+}$.
@@
is increasing because $(\theta^{2}-\sin^{2}\theta)^{\prime}=2\theta-\sin(2\theta)>0$. So the $\max\phi(\theta)$ over $[0,\mu_{-}]$ is attained at $\theta=\mu_{-}$.
@@
Since $01\geq\mu=\sin\theta_{k}>\sin\theta>$ 0 and $\theta>0$, we have
@@
$\phi(\theta)=\frac{\theta}{\sin^{2}\theta-\mu^{2}}-\frac{\theta}{\theta^{2}-\mu^{2}}=\frac{1}{2\mu}\left(\frac{\theta}{\sin\theta-\mu}-\frac{\theta}{\sin\theta+\mu}\right)-\frac{1}{2}\left(\frac{1}{\theta-\mu}+\frac{1}{\theta+\mu}\right).$
We will show that $\phi(\theta)$ is convex for $\theta\in(\theta_{k},\frac{\pi}{2})$. We calculate
$\phi^{\prime\prime}(\theta)=$ $\frac{\theta\sin\theta-2\cos\theta}{2\mu}\left(\frac{1}{(\sin\theta-\mu)^{2}}-\frac{1}{(\sin\theta+\mu)^{2}}\right)+\frac{\theta\cos^{2}\theta}{\mu}\left(\frac{1}{(\sin\theta-\mu)^{3}}-\frac{1}{(\sin\theta+\mu)^{3}}\right)$
$-\frac{1}{(\theta-\mu)^{3}}-\frac{1}{(\theta+\mu)^{3}}\geq F(s,\theta,\mu):=$
$\frac{\theta s-1}{\mu}\left(\frac{1}{(s-\mu)^{2}}-\frac{1}{(s+\mu)^{2}}\right)+\frac{\theta(1-s^{2})}{\mu}\left(\frac{1}{(s-\mu)^{3}}-\frac{1}{(s+\mu)^{3}}\right)-\frac{1}{(\theta-\mu)^{3}}-\frac{1}{(\theta+\mu)^{3}},$
where we used $\cos\theta<1-\frac{1}{2}\theta\sin\theta$ for $0<\theta<\frac{\pi}{2}$ and $s:=\sin\theta$. After multiplying $F(s,\theta,\mu)$ with $(\theta-\mu)^{3}(\theta+\mu)^{3}(s-\mu)^{3}(s+\mu)^{3}$, we get a polynomial in $\mu^{2}$ with degree at most four. Then we can verify that the polynomial is always positive by the software Mathematica as follows
  Resolve[Exists[{s,t,v}, p<=0 && s>t-t^3/6 && s<t-t^3/6+t^5/120 && s<1 && 0<t
                 && t<Pi/2 && 0<v && v<=1], Reals]
where $p$ stands for the polynomial, $t$ for $\theta$ and $v$ for $\mu^{2}$. We used the inequalities $\theta-\frac{\theta^{3}}{6}<\sin\theta<\theta-\frac{\theta^{3}}{6}+\frac{\theta^{5}}{120}$ for $\theta\in(0,\frac{\pi}{2})$, and the assumption $10<\mu<$ 1. To the above command, Mathematica gives the answer false, which means the polynomial is always positive. Hence, $\phi(\theta)$ on $\theta\in[\theta_{+},\theta_{\max}]$ attains its maximum at either $\theta_{+}$ or $\theta_{\max}$ when $\theta_{+}<\theta_{\max}$ (otherwise, there is no such interval to consider). Note that if $\mu\leq C_{\mu}<\frac{1}{2}$ and $k>\pi$ then it holds that $\theta_{+}<\theta_{\max}$. Indeed, we have $h<\frac{1}{\pi}<\frac{2}{\pi+\sigma_{+}^{h}}$ so $\sin(\theta_{\max}-\frac{\sigma_{+}^{h}h}{2})=\cos(\frac{\pi+\sigma_{+}^{h}}{2}h)>1-\frac{(\pi+\sigma_{+}^{h})^{2}}{8}h^{2}>\frac{1}{2}>\sin\theta_{k}$.
@@
Note that $\theta_{k}>\theta_{-}\geq\mu_{+}>\mu$ and $\arcsin x-x<x^{3}$ for $10<x<$ 1. So
@@
[…]
@@
$\frac{1}{2\mu_{+}}<\frac{1}{\mu+\mu_{+}}<\frac{1}{2\mu}\quad\text{ instead of }\quad\frac{1}{2}<\frac{\mu_{+}}{\mu+\mu_{+}}<1$
@@
When $\theta_{k}<\theta<\frac{\pi}{2}$, with $\mu=\sin\theta_{k}$ we have
$\phi_{e}(\theta)=\frac{1}{\sin^{2}\theta-\mu^{2}}-\frac{1}{\theta^{2}-\mu^{2}}.$
It can be verified by the software Mathematica that $\phi_{e}$ is convex given $\theta\in(\theta_{k},\frac{\pi}{2})$ and $10<\mu<$ 1; see the proof of Lemma 10 for more details of the approach. So the candidates for $\max\phi_{e}$ on $\Theta\cap(\theta_{k},\theta_{\max}]$ are $\phi_{e}(\theta_{+})$ (if $\theta_{+}<\theta_{\max}$) and $\phi_{e}(\theta_{\max})$. To estimate $\phi_{e}(\theta_{+})$, we can again modify the estimates of $\phi(\theta_{+})$ in the proof of Lemma 10. Using
$\frac{1}{2\theta_{+}}<\frac{1}{\mu+\theta_{+}}<\frac{1}{2\mu}\quad\text{ instead of }\quad\frac{1}{2}<\frac{\theta_{+}}{\mu+\theta_{+}}<1,$
@@
Since $\theta_{\max}^{2}=\frac{\pi^{2}}{4}(1-2h+h^{2})$, $10<\mu<$ 1 and $0<h\leq\frac{1}{4}$, we find
@@
Denote $\theta_{e}:={\xi_{e}h}/{2}$. To find the minimal point $\theta_{e}$ of $\phi_{e}$ on $[\theta_{k},\frac{\pi}{2}]$, we calculate
@@
Since $\phi_{e}$ is convex in $[\theta_{k},\frac{\pi}{2}]$ and $\theta_{e}=\arg\min_{[\theta_{k},{\pi}/{2}]}\phi_{e}$, we have $\phi_{e}^{\prime}(\theta_{e})=0$, from which we find
@@
Substituting this into $\phi_{e}(\theta_{e})$, and using $\frac{x}{\sin x}\!>\!1\!+\!\frac{x^{2}}{6}$ for $0\!<\!x\!<\!\pi$ and $\sqrt{1+x}\!>\!1\!+\!\frac{x}{3}$ for $30\!<\!x\!<\!$ 3, we get
@@
$\tilde{\sigma}_{k}:=\min_{\xi,h}\left|k-\frac{2}{h}\sin\frac{\xi h}{2}\right|>0\;\text{over $\xi\in\{1,\ldots,N-1\}\pi$ and the sequence of $h$ going to zero}.$
@@
and when $k^{3}h^{2}>C_{2}$ it holds that $C(k,h)$ or $kC_{e}(k,h)$ is greater than a positive constant. Here $C_{e}(k,h)$ is multiplied with $k$ because $|E_{1}^{h}|_{1}\leq C_{e}(k,h)|f_{\mathrm{low}}|_{1}$ holds with equality when $f_{\mathrm{low}}=\sin(\xi x)$ for some $\xi\in\pi\mathbb{N}$ satisfying $|\xi-k|<\pi$ and in this case $|f_{\mathrm{low}}|_{1}$ is of order $k$.

### v2

@@ 3.3 Aliasing error of the classical 3-point centred scheme
$=\sum_{(\xi/\pi)=1}^{N-1}\left(\frac{\xi}{\lambda}-\frac{\xi}{\lambda^{h}}\right)\hat{f}(\xi)\cos(\xi x)-\sum_{(\xi/\pi)=1}^{N-1}\frac{\xi}{\lambda^{h}}\sum_{s=\pm 1}s\sum_{m=1}^{\infty}\hat{f}(s\xi+2mN\pi)\cos(\xi x).$
@@
Suppose $k>\pi$, $0<\frac{kh}{2}\leq C_{\mu}<1$, and there is a sequence of $h=\frac{1}{N}$ going to zero such that
[…]
@@
because by the mean value theorem $\tilde{\theta}<\theta_{+}$. Note that the derivative of the numerator of $\phi^{\prime}(\theta)$ is $2\theta\cos(2\theta)$. So if $\theta_{+}<\frac{\pi}{4}$, then $\phi^{\prime}<0$ on $\theta_{+},\frac{\pi}{4}]$. Therefore, no matter $\theta_{+}<\frac{\pi}{4}$ or not, any stationary point $\theta_{c}$ of $\phi$ in $\theta_{+},\theta_{\max}]$ must satisfy $\theta_{c}\geq\frac{\pi}{4}$. It follows that
@@
By Lemma 8 the second part $(E_{2}^{h})^{\prime}$ of $(e_{1}^{h})^{\prime}$ in (22) is controlled, similarly to $(e_{2}^{h})^{\prime}$ in (4), by the grid under-resolved source $f_{\mathrm{high}}$, and behaves as $O(h^{p})$ when $k>\pi$, $\frac{kh}{2}\leq C_{\mu}<1$ and $\tilde{\sigma}_{k}>0$.
@@
Suppose $k>\pi$, $0<\frac{kh}{2}\leq C_{\mu}<1$, and there is a sequence of $h=\frac{1}{N}$ going to zero such that
[…]
@@
$\leq$  $2\max_{\xi\in\{1,..N-1\}\pi}\frac{1}{|\lambda^{h}|^{2}}\sum_{(\xi/\pi)=1}^{N-1}\left[\left|\sum_{m=1}^{\infty}\hat{f}(\xi+2mN\pi)\right|^{2}+\left|\sum_{m=1}^{\infty}\hat{f}(-\xi+2mN\pi)\right|^{2}\right].$
@@ 3.4 Absolute error estimates of the classical 3-point centred scheme
[…]
@@
where $\psi$ is given in (30), $\psi(k_{\pm}^{h})$ can be removed if $k_{\pm}^{h}\geq\xi_{\max}$, and $k_{+}$ (resp. $k_{-}^{h}$) equals $k_{+}^{h}$ (resp. $k_{-}$) when $\pi\mathbb{N}\cap(k,k^{h})=\emptyset$. Moreover, we have
@@
$\frac{1}{15}\frac{(k+\sigma_{+})^{3}h^{2}}{\sigma_{+}(4\pi+k^{3}h^{2})}<$ $\,\psi(k_{+})<\frac{2}{3}\frac{k^{3}h^{2}}{\sigma_{-}^{h}\sigma_{+}\sqrt{1-C_{\mu}^{2}}}\,$, $\text{ if }\pi\mathbb{N}\cap(k,k^{h})\neq\emptyset,$
$\frac{(1-\frac{\pi^{2}}{80}C_{\mu}^{2})(k+\sigma_{+})^{3}h^{2}}{12\pi(4\sigma_{+}+k^{3}h^{2})}<$ $\,\psi(k_{-}^{h})<\frac{\pi^{4}}{384}\frac{k^{3}h^{2}}{\sigma_{-}^{h}\sigma_{+}\sqrt{1-C_{\mu}^{2}}}\,$, $\text{ if }\pi\mathbb{N}\cap(k,k^{h})\neq\emptyset,$
$\frac{(1-\frac{(1+\pi)^{2}}{80}C_{\mu}^{2})(k+\sigma_{+}^{h})^{3}h^{2}}{12\pi(4\sigma_{+}^{h}+k^{3}h^{2})}<\,$ $\psi(k_{+}^{h})<\frac{(1+\pi)^{4}}{16}\frac{k^{3}h^{2}}{\sigma_{+}^{h}(24\sigma_{+}^{h}+k^{3}h^{2})\cos(\theta_{\mu}+\frac{\sigma_{+}^{h}h}{2})}\,$, $\text{ if }C_{\mu}<\cos\frac{\sigma_{+}^{h}h}{2},$
$\frac{h}{\pi}\left(\frac{9\pi^{2}}{64}\!-\!1\right)<$ $\,\psi(\xi_{\max})<\frac{4(\pi^{2}-4)\pi h}{\left(\cos^{2}\frac{\pi h}{2}\!-\!C_{\mu}^{2}\right)(9\pi^{2}-64)}\,$, $\text{ if }C_{\mu}<\cos\frac{\pi h}{2},$
@@
Note that $\pi\mathbb{N}\cap(k,k^{h})\neq\emptyset$ if and only if $k_{+}<k^{h}$. Using $\mu^{3}/6<\arcsin\mu-\mu<(\pi-2)\mu^{3}/2$ for $0<\mu<1$, we find a sufficient condition for $k_{+}\geq k^{h}$ is that $k^{3}h^{2}\leq 8\sigma_{+}/(\pi-2)$ and a sufficient condition for $k_{+}<k^{h}$ is that $k^{3}h^{2}>24\sigma_{+}$.
@@
is increasing because the numerator is increasing and the denominator is decreasing. Indeed, $(\theta^{2}-\sin^{2}\theta)^{\prime}=2\theta-\sin(2\theta)>0$. So the $\max\phi(\theta)$ over $0,\mu_{-}]$ is attained at $\theta=\mu_{-}$.
@@
Since $1\geq\mu=\sin\theta_{k}>\sin\theta>0$ and $\theta>0$, we have
@@
[…]
The proof can be found in Appendix A. Hence, $\phi(\theta)$ on $\theta\in[\theta_{+},\theta_{\max}]$ attains its maximum at either $\theta_{+}$ or $\theta_{\max}$ when $\theta_{+}<\theta_{\max}$ (otherwise, there is no such interval to consider). Note that if $\mu\leq C_{\mu}<\frac{1}{2}$ and $k>\pi$ then it holds that $\theta_{+}<\theta_{\max}$. Indeed, we have $h<\frac{1}{\pi}<\frac{2}{\pi+\sigma_{+}^{h}}$ so $\sin(\theta_{\max}-\frac{\sigma_{+}^{h}h}{2})=\cos(\frac{\pi+\sigma_{+}^{h}}{2}h)>1-\frac{(\pi+\sigma_{+}^{h})^{2}}{8}h^{2}>\frac{1}{2}>\sin\theta_{k}$.
@@
Note that $\theta_{k}>\theta_{-}\geq\mu_{+}>\mu$ and $\arcsin x-x<x^{3}$ for $0<x<1$. So
@@
$\frac{1}{15}\frac{(k+\sigma_{+})^{2}h^{2}}{\sigma_{+}(4\pi+k^{3}h^{2})}<\,\psi_{e}(k_{+})$ $\,<\frac{1}{3}\frac{k^{2}h^{2}}{\sigma_{-}^{h}\sigma_{+}\sqrt{1-C_{\mu}^{2}}},\,$ $\text{ if }\pi\mathbb{N}\cap(k,k^{h})\neq\emptyset,$
$\frac{(1-\frac{\pi^{2}}{80}C_{\mu}^{2})(k+\sigma_{+})^{2}h^{2}}{3\pi(2+\pi)(4\sigma_{+}+k^{3}h^{2})}<\,\psi_{e}(k_{-}^{h})\,$ $<\frac{\pi^{4}}{768}\frac{k^{2}h^{2}}{\sigma_{-}^{h}\sigma_{+}\sqrt{1-C_{\mu}^{2}}},\,$ $\text{ if }\pi\mathbb{N}\cap(k,k^{h})\neq\emptyset,$
$\frac{(1-\frac{(1+\pi)^{2}}{80}C_{\mu}^{2})(k+\sigma_{+}^{h})^{2}h^{2}}{12\pi(4\sigma_{+}^{h}+k^{3}h^{2})}<\,\psi_{e}(k_{+}^{h})\,$ $<\frac{(1+\pi)^{4}k^{2}h^{2}}{32\sigma_{+}^{h}(24\sigma_{+}^{h}\!+\!k^{3}h^{2})\cos(\theta_{\mu}\!+\!\frac{\sigma_{+}^{h}h}{2})},\,$ $\text{ if }C_{\mu}<\cos\frac{\sigma_{+}^{h}h}{2},$
$\frac{(9\pi^{2}-64)h^{2}}{64\pi^{2}}<\psi_{e}(\xi_{\max})$ $\,<\frac{4(\pi^{2}-4)h^{2}}{\left(\cos^{2}\frac{\pi h}{2}-C_{\mu}^{2}\right)\left(9\pi^{2}-64\right)},\,$ $\text{ if }C_{\mu}<\cos\frac{\pi h}{2},$
@@
[…]
@@
When $\theta_{k}<\theta<\frac{\pi}{2}$, with $\mu=\sin\theta_{k}$ we have that
[…]
The proof can be found in Appendix B. So the candidates for $\max\phi_{e}$ on $\Theta\cap(\theta_{k},\theta_{\max}]$ are $\phi_{e}(\theta_{+})$ (if $\theta_{+}<\theta_{\max}$) and $\phi_{e}(\theta_{\max})$. To estimate $\phi_{e}(\theta_{+})$, we can again modify the estimates of $\phi(\theta_{+})$ in the proof of Lemma 10. Using
[…]
@@
Since $\theta_{\max}^{2}=\frac{\pi^{2}}{4}(1-2h+h^{2})$, $0<\mu<1$ and $0<h\leq\frac{1}{4}$, we find
@@
Denote $\theta_{e}:={\xi_{e}h}/{2}$. To find the minimal point $\theta_{e}$ of $\phi_{e}$ on $\theta_{k},\frac{\pi}{2}]$, we calculate
@@
Since $\phi_{e}$ is convex in $\theta_{k},\frac{\pi}{2}]$ and $\theta_{e}=\arg\min_{[\theta_{k},{\pi}/{2}]}\phi_{e}$, we have $\phi_{e}^{\prime}(\theta_{e})=0$, from which we find
@@
Substituting this into $\phi_{e}(\theta_{e})$, and using $\frac{x}{\sin x}\!>\!1\!+\!\frac{x^{2}}{6}$ for $0\!<\!x\!<\!\pi$ and $\sqrt{1+x}\!>\!1\!+\!\frac{x}{3}$ for 0<x<30\!<\!x\!<\!3, we get
@@
[…]
@@
and when $k^{3}h^{2}>C_{2}$ it holds that $C(k,h)$ or $kC_{e}(k,h)$ is greater than a positive constant.
Remark 11.
Here $C_{e}(k,h)$ is multiplied with $k$ because $|E_{1}^{h}|_{1}\leq C_{e}(k,h)|f_{\mathrm{low}}|_{1}$ holds with equality when $f_{\mathrm{low}}=\sin(\xi x)$ for some $\xi\in\pi\mathbb{N}$ satisfying $|\xi-k|<\pi$ and in this case $|f_{\mathrm{low}}|_{1}$ is of order $k$. Note that the upper bound is attained when $f$ is chosen dependently on $k$ and as a result $|f|_{p}=O(k^{p})$.